Рelen found a small golden colored bead with a volume of 0.50mL. A) in order for the bead to be pure gold what mass will it have? B) helen found the mass fo the bead to be 3.65g; did she find a gold bead? if not, what might it be?


1
Expert's answer
2020-02-12T05:48:34-0500

A. To identify the mass of a golden bead, volume and density values must be used:

m = d × V,

where m - mass, d - density, V - volume.

As the density of gold is 19.3 g/cm3 (=19.3 g/ml):

m = 19.3 g/ml × 0.50 ml = 9.65 g.

B. As the mass of the bead is 3.65 g, the bead is not made of gold. The real density of the bead is:

d = m / V = 3.65 g / 0.50ml = 7.3 g/ml = 7.3 g/cm3

The bead may be made of a tin (density 7.26 g/cm3).